Pearson Motors has a target capital structure of 35% debt and 65% common equity, with no preferred stock. The yield to maturity on the company's outstanding bonds is 10%, and its tax rate is 25%. Pearson's CFO estimates that the company's WACC is 11.30%. What is Pearson's cost of common equity? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1

WACC=(Weight of equity)*(Cost of equity)+(Weight of debt)*(Cost of debt)*(1-tax)
11.30%=(65%)*(Cost of equity)+(35%)*(10%)*(1-25%)

11.30=(65)*(Cost of equity)+(35)*0.075
11.30=(65)*(Cost of equity)+2.625
11.30-2.625=(65)*(Cost of equity)
8.675=(65)*(Cost of equity)

Cost of equity=8.675/65=0.133461 or 13.35% (Rounded to two decimal places)